Bill and Mary plan to marry in December 2018. Bill​'s salary is $105,000 and he owns his residence. His itemized deductions total $ 19 comma 000. Mary​'s salary is $86,000. Her itemized deductions total only $8,100 as she does not own her residence. For purposes of this​ problem, assume 2019 tax rates and standard deductions are the same as 2018.

Requirements:

A. What will their 2018 tax be if they marry before year-end and file a joint return?

adjusted Gross income ?
minus:
itemized deductions ?
taxable income ?

B. What will their combined 2018 taxes be if they delay the marriage until 2019?

C. What factors contribute to the difference in taxes?
Answer #1

Part A

Adjusted gross income (105000+86000) 191000
Deductions from AGI
Greater of standard deductions or itemized deductions (24000 or (19000+8100)) (27100)
Taxable income 163900
Tax liability 27937

Tax 2018(married filling jointly ) = 8907+(22%*(163900-77400)) = 27937

Part B

Bill's tax

AGI 105000
Deductions from AGI
Greater of standard deductions or itemized deductions (12000 or 19000) (19000)
Taxable income 86000
Bill's tax 14089.50+(24%*(86000-82500)) 14929.50

Mary's tax

AGI 86000
Deductions from AGI
Greater of standard deductions or itemized deductions (12000 or 8100) (12000)
Taxable income 74000
Mary's tax 4453.50+(22%*(74000-38700)) 12219.50

Combined tax = 14929.50+12219.50 = $27149

Part C

Mainly, it is the filling status which contributes significantly to the difference in taxes. Moreover, the type of property, gambling, medical expenses, etc. affects the tax amount.